The equation of a common tangant to [x^2/a^2] - [y^2/b^2]=1 and [y^2/a^2] - [x^2/b^2]=1 is (1)y=x+{a^2-b^2}^1/2 (2)y=x-{a^2-b^2}^1/2 (3)y=-x+{a^2-b^2}^1/2 (4)y=-x-{a^2-b^2}^1/2

Grade:

1 Answers

vikas askiitian expert
13 years ago

for hyperbola , x2/a2 - y2/b2 = 1

y intercept of tangent of hyperbola = c = (a2m2- b2)1/2

 

for hyperbola , y2/b2 - x2/a2

y intersept of tangent of hyperbola = c = (b2m2- a2)1/2 

for same line , intercept are equal so

a2m2-b2 = b2m2 - a2

m = -1 so

c =

eq of tangent = y = mx + c

option 3,4 are correct